OpenStudy (anonymous):

In 4x^2-25y^2z^2, would the answer be (2x+5yz)(2x-5yz)?

OpenStudy (anonymous):

I am not sure on this one. It was harder to solve. Could you please double check my work?

OpenStudy (anonymous):

what is the ln? is it \[\ln(4x^2-25x^z^2)\]

OpenStudy (anonymous):

i mean \[\ln(4x^2-25y^2z^2)\]

OpenStudy (anonymous):

or is that 'ln' just a typo?

OpenStudy (anonymous):

if it is \[\ln(4x^2-25y^2z^2)\] then you factored properly to get \[\ln((2x+5yz)(2x+5yz))=\ln(2x+5yz)+\ln(2x-5yz)\]

OpenStudy (anonymous):

there is no ln in the problem. :(

OpenStudy (anonymous):

no that is In not LN.

OpenStudy (anonymous):

I think he was saying 'in' not 'ln', right?!

OpenStudy (anonymous):

then you were right at the beginning and i should just be quiet